A net force of 40 N acts on a 20 kg object. Find the object's acceleration.
1
Expert's answer
2020-12-06T17:25:33-0500

According to Newton's second law of motion, the acceleration of an object equals the net force acting on it divided by its mass, or a = "\frac{F}{m}" .


F= 40 N, m= 20 kg.


a = "\frac{40}{20}" = 2 (m/s2).